sinA/cotA+cosecA
1
Expert's answer
2011-05-24T08:37:59-0400
sinA/ cotA + cosec A = sinA/(cosA/sinA) + 1/sinA = sin2A/cos A +1/sinA = (sin3A + cos A)/(sinA cosA)